# {Authlist}

### v1

@@
[…]
{Authlist}
C. Amsler1, H. Breuker2, M. Bumbar3,4, M. Cerwenka1,4, G. Costantini5,6, R. Ferragut7,8, M. Fleck1††thanks: Present address: University College London, Department of Physics & Astronomy, London, United Kingdom., M. Giammarchi8, A. Gligorova1††thanks: Present address: University of Vienna, Faculty of Physics, Vienna, Austria.††thanks: Electronic mail: angela.gligorova@oeaw.ac.at, G. Gosta5,6, E. D. Hunter1††thanks: Present address: Experimental Physics Department, CERN, Geneva, Switzerland.,, C. Killian1, B. Kolbinger1††thanks: Present address: DATA TECHNOLOGY, Vienna, Austria., V. Kraxberger1,4, N. Kuroda9, M. Lackner1, M. Leali5,6, G. Maero8,10, V. Mascagna5,6, Y. Matsuda9, S. Migliorati5,6, D. J. Murtagh1, A. Nanda1,4, L. Nowak3,4, S. Rheinfrank1, M. Romé8,10, M. C. Simon1, M. Tajima11††thanks: Present address: Japan Synchrotron Radiation Research Institute, Hyogo, Japan., V. Toso8,10, S. Ulmer2, M. van Beuzekom12, L. Venturelli5,6, A. Weiser1,4, E. Widmann1, Y. Yamazaki2, J. Zmeskal1
[…]
1Stefan Meyer Institute for Subatomic Physics, Austrian Academy of Sciences, Vienna, Austria
2Ulmer Fundamental Symmetries Laboratory, RIKEN, Saitama, Japan
3Experimental Physics Department, CERN, Geneva, Switzerland
4University of Vienna, Vienna Doctoral School in Physics, Vienna, Austria
5Dipartimento di Ingegneria dell’Informazione, Università degli Studi di Brescia, Brescia, Italy
6INFN sez. Pavia, Pavia, Italy
7L-NESS and Department of Physics, Politecnico di Milano, Como, Italy
8INFN sez. Milano, Milan, Italy
9Institute of Physics, Graduate School of Arts and Sciences, University of Tokyo, Tokyo, Japan
10Dipartimento di Fisica, Università degli Studi di Milano, Milan, Italy
11RIKEN Nishina Center for Accelerator-Based Science, Saitama, Japan
12Nikhef National Institute for Subatomic Physics, Amsterdam, Netherlands
@@
The mechanism of antiproton-nucleus annihilation at rest is not fully understood, despite substantial previous experimental and theoretical work. In this study we used slow extracted, sub-keV antiprotons from the ASACUSA apparatus at CERN to measure the charged particle multiplicities and their energy deposits from antiproton annihilations at rest on three different nuclei: carbon, molybdenum and gold. The results are compared with predictions from different models in the simulation tools Geant4 and FLUKA. A model that accounts for all the observed features is still missing, as well as measurements at low energies, to validate such models.
@@
Keywords: antiproton annihilation, nucleus, Monte Carlo simulations, annihilation products, low energy, nuclear reaction
1 Introduction
@@
Antiproton () annihilation on a nucleus (A) is a particular nuclear reaction in which a relatively large amount of energy is released. Part of this energy is transferred to the nucleus with reduced linear and angular momentum, compared to collisions with protons or heavy ions. Consequently, the decay mechanism of nuclei heated by stopped and energetic antiprotons allows the observation of some effects characteristic of heavy-ion reactions, such as multifragment disintegration, without nuclear compression [1, 2, 3]. Our understanding of the mechanism of antiproton annihilation in nuclear matter is based on observables such as multiplicity, energy and angular distributions of secondary particles. Various models, including the Intranuclear Cascade (INC) model [4, 5], the statistical multifragmentation model [6] or the Lanzhou quantum molecular-dynamics approach [7] describe the nuclear dynamics and decay mechanism induced by antiprotons with varying degrees of success. A great amount of complexity in the antiproton-nucleus annihilation comes from the rich variety of possible reaction channels triggered by the primary annihilation mesons.
@@
Annihilation at rest most of the time occurs on a single nucleon in the peripheral nuclear region, where the nucleon density is less than 10% of the central density [5]. For annihilation, the total energy of 1877 MeV is converted into an average of 3.00.2 charged pions () and 2.00.2 neutral pions (), with an average kinetic energy of 230 MeV [8, 9]. Kaons and mesons are produced in about 6% and 7% of all annihilations, respectively. When the annihilation occurs on a neutron, on average 1.070.04 and 2.070.04 are produced [10]. Absorption or scattering (final state interactions, FSI) of the emitted mesons may follow, with or without break-up of the nucleus, which depends upon the solid angle under which the remaining nucleons are seen from the annihilation point [11], and which may modify the detected meson distribution [12]. While many annihilation channels have been measured for antiproton-nucleon (N) annihilation at rest, for A the branching ratios have been hitherto reported only for and [13, 10, 14], and the total prong multiplicity has been measured for a very limited number of nuclei [15]. Furthermore, even for annihilation the current models do not perform satisfactorily in describing all the aspects of the measured data.
@@
The bulk of data originates from LEAR (Low Energy Antiproton Ring) experiments, taken about three decades ago by degrading 20 MeV antiprotons in a number of moderators, and ultimately stopping them in thick targets. In these measurements only light prongs (charged particles) were detected (helium and lithium ions being the heaviest), in certain energy intervals and within narrow solid angles [3, 16, 1, 17]. Features such as particle yields and momentum spectra of the emitted pions, protons (p), deuterons (d) and tritons (t) with energies above 10 MeV were reported, as well as the energy dissipation (energy transfer to the nucleons and mean excitation energies [3]). The residual nuclei have been investigated through measurements of their radioactivity after irradiation, i.e. of their gamma spectra [18, 19, 20, 21, 22, 23], and the neutron density distributions were deduced through nuclear spectroscopy analysis of the antiproton annihilation residues [24]. The experimental results have been most extensively compared to the INC, which models the A annihilation through the antiproton annihilation on a single nucleon, generating a number of pions which then cascade through the nucleus in a sequence of interactions with the rest of the nucleons. Good agreement with data was found on the absolute pion and proton yields, however the momentum distribution of pions and protons emerging from the annihilation were described with less accuracy [25]. For the residue distribution, the evaporation model provides a satisfactory description [26].
@@
Besides its relevance for nuclear physics studies, the antiproton-nucleus annihilation at rest is one of the key processes for experiments at the Antiproton Decelerator (AD) at CERN, that study antimatter in laboratory, detecting it through annihilation [27, 28, 29, 30, 31, 32]. Monte Carlo simulations are performed to assess the efficiency of tagging antihydrogen events, relying on the physics assumptions implemented either in Geant4, such as the CHiral Invariant Phase Space event generator (CHIPS) [33, 34, 35], Fritiof-Precompound (FTFP) [36, 37, 38] and the newly added Intranuclear Cascade de Liège (INCL) model [39], or in FLUKA [40, 41, 42]. The low energy annihilation models used in these software packages are, however, based on hadronic high energy interactions (CHIPS, FTFP) or were developed for medical physics applications (FLUKA), and were extrapolated to low energies in spite of the fact that the low-energy annihilation mechanism is still not well understood [43]. Recent studies have been performed using 5 MeV pulsed antiproton beam from the AD, degrading it to energies of 100 keV in a broad distribution. Antiprotons annihilating at rest in Al, Cu, Ag, Au and Si were measured with emulsion and silicon detectors, as reported in [44, 45, 46]. The results revealed significant discrepancies in the prong multiplicity between measured data and the Geant4 models, with differences ranging from 30% to a factor of 4.
In this work, we present new measurements of A annihilation at rest using a slow extracted, monoenergetic beam of 150 eV, accelerated to 1.15 keV by the electric potential applied on the target. The antiprotons annihilated on thin targets, providing a crucial comparison with the models in which the stopping effects and production of secondaries due to the particle propagation can be neglected. In addition, this allows the emergence and detection of the heavy nuclei produced in the annihilation.
@@ 2 Experimental setup
The antiprotons were annihilated on three different targets: 2 thick diamond-like-carbon (DLC) and molybdenum foils, and a gold foil of 1 thickness. The stopping range of 1.15 keV antiprotons, according to ref. [53], is 46 nm for carbon and 57 nm for gold. The beam was centered and focused on the target with an Einzel lens (see fig. 1), where both the entrance and exit electrodes have additional diagonal cuts. Consisting of five electrodes, this lens allows for two-dimensional steering perpendicularly to the beam axis.
The on-target beam size was determined by first measuring the antiproton beam on the bismuth germanium oxide (BGO) crystal detector used for the antihydrogen experiment [54], and by using the results of 30 mm FWHM in x and y direction [55] as an input to simulations. As the foils then replaced the BGO for the current measurements, the beam size on the target foils was estimated from the measured size of the beam, adding, in the Geant4 simulations, the 1 kV electric potential applied on the foil for further focusing. The outcome of this combined work is shown in Fig. LABEL:fig:XY_pbar_beam_trig_timea and LABEL:fig:XY_pbar_beam_trig_timeb, where the x-y beam profile and its horizontal (x) and vertical (y) projections are shown, respectively. A beam size of 10.5 mm in x and y (FWHM) was obtained, with 92.4 of the extracted antiprotons annihilating on the 22 target. These results were also confirmed with SIMION [56] simulations, starting from the same initial conditions and including the voltage bias on the foil.
@@ 2.1 Detection and identification of charged pions and heavy fragments
The charged annihilation products were simultaneously detected by two detectors, a cylindrical hodoscope surrounding the target and a pixel detector placed 1 cm behind it (Timepix3 quad), as shown in Fig. 3. Pions and heavier fragments (p, d, t, He, Li, etc.) in the forward direction were impinging on the 22 array of Timepix3, with a total area of 2.82.8 . The ASICs (application-specific integrated circuit), containing 256256 pixel channels of 5555 each, were coupled to a 500 thick silicon sensor for this application. Providing a simultaneous information of the time-of-arrival (ToA) and the time-over-threshold (ToT) for each pixel, with a nominal time resolution of 1.56 ns, the detector is able to detect mixed radiation fields and has a long history of a variety of applications [57, 58, 59, 60, 61]. Its wide dynamic spectral range allows for detection of all charged particles depositing energy above the threshold. In this experiment, the detection threshold was set to of collected charge per pixel, corresponding to a minimum required energy of 3.6 keV for X rays and low-energy gamma rays, tens of keV for electrons, hundreds of keV for protons and MeV for heavy ions [62]. The higher threshold for heavier particles originates from the aluminum electrode, approximately 500 nm in thickness, that is deposited on top of the silicon sensor for biasing purposes. For details about the ASIC and the readout system see ref. [63, 64, 65].
@@
The two detectors were time synchronised, and the data acquisition coincided with the start of the slow extraction cycle. Every time a hit was detected in both inner and outer bar layers of the hodoscope, a trigger was issued and this event was time-stamped in the Timepix3 data stream, where every hit in this detector was also recorded. The time distribution of these events (triggers) for one extraction cycle is shown in Fig. LABEL:fig:triggers_hitsa, revealing the 20 s duration of the antiproton beam extracted from MUSASHI within the 140 s long cycle. About 1,000 annihilation events per slow extraction cycle were detected, resulting in a total of 100,000 annihilation events for each target foil. The triggers outside the slow extraction window are produced by cosmic rays or, sporadically, from secondaries of other origin in the AD experimental hall crossing the hodoscope. The two small peaks in the Timepix3 distribution are from two cosmic particles that traversed both the hodoscope and the Timepix3 detectors. The data analysis was performed only within the slow extraction time window.
@@
The hits in the Timepix3 were assigned to a particular event (trigger) according to their ToA. The interval between two successive triggers ranged from 10 ms to a few 10 ms, as illustrated in Fig. LABEL:fig:triggers_hitsb. The observed cutoff at 10 ms stems from the hodoscope data acquisition limit of 100 Hz. The nominal cosmic background was measured to be 1.3 Hz. When combined with the duration of one extraction (20 s) it gives a total of 26 cosmics events, which is only 2.6% of the average number of annihilation events recorded per run. For the Timepix3 the cosmic ray background is fully negligible.
@@ 2.1.2 Charged pions and heavier fragments in Timepix3
In the pixel detector, charged particles were detected in 25% of the full solid angle. Based on the ToT proportionality to the amount of charge collected in each electrode, a per pixel energy calibration was carried out using test pulses. The deposited energy was extracted from the known value of the test pulse capacitor, assuming an energy of 3.6 eV for the creation of one electron-hole pair in silicon. The calibration results were validated with data from a radioactive source and cosmic rays. In Fig. LABEL:fig:Am241_energya, the lower end of the spectrum, converted to energy units, is presented with a Gaussian fit to the characteristic 59.5 keV peak, determined to be 59.32.6 keV.
@@
The energy deposit and morphology of the pixelated track a charged particle generates in the Timepix3 (also referred to as pixel cluster) are influenced by factors such as the particle type, its kinetic energy, and the incidence angle. The graph in Fig. LABEL:fig:Am241_energyb illustrates the distribution for cosmic rays, both in measurements and simulations, exclusively for clusters with high (0.8) eccentricity, a measure of the linearity of the shape of a particle track. These elongated tracks correspond to cosmic rays that cross the Timepix3 at large angles with respect to its normal. The distribution is modeled with a Landau-Gaussian fit and compared to Geant4 simulations. The most probable value of 1.4 agrees with the typical value characteristic for MIPs in silicon [68].
@@
Due to the geometry of the set-up (see Fig. 3), the annihilation prongs hit the pixel detector only at acute angles with respect to its normal. The maximum angle is 55 for prongs emerging from the foil’s center, and up to 67 for prongs originating from annihilations close to the edges of the target. Most of the particles impinge on the Timepix3 at angles smaller than these, which results in the formation of mostly blob-like clusters and relatively short tracks, with a maximum length of less than 1.5 mm. Combined with the variety of particles produced in the annihilation and their diverse ranges of energies and incident angles, this allowed only a broad classification of heavily ionizing particles (HIPs) as opposed to MIPs. Ongoing efforts from a number of users of Timepix ASICs are aimed at developing tools for more precise particle identification in mixed radiation fields [62].
@@
In this study, the differentiation of MIPs in the Timepix3 was achieved by first implementing cuts based on the energy deposit in a cluster. These cuts were derived from the analysis of cosmic data. The energy thresholds for the highest pixel energy () within the cluster and the mean pixel energy () were extracted from the versus distribution depicted in Fig. LABEL:fig:cosmics_cutsa, by calculating its cumulative distribution function (CDF), shown in Fig. LABEL:fig:cosmics_cutsb. The threshold was set to account for 99 of the cosmic rays data sample. The remaining clusters underwent further evaluation based on their morphology and were categorized as small blobs, large blobs, or heavy tracks before being identified as HIPs. The efficiency of removing MIPs from the annihilation clusters using only energy cuts from cosmics was estimated, from simulations, to be 97. However, it was also revealed that some high-energy protons can be mistakenly identified as MIPs after applying these cuts, as shown in Fig. LABEL:fig:Etot_MIPs_tag. The comparison plots show the deposited energy by charged pions (solid line) and by MIPs tagged after applying the energy cuts obtained from simulated cosmic rays (dashed lines) for two Geant4 models ((a) FTF and (b) CHIPS), for all three targets. Additionally, the difference between the two energy depositions is presented at the bottom. The excess counts are attributed to the aforementioned protons, the quantity of which depends on the specific simulation model. The percentage of these protons incorrectly tagged as MIPs relative to the total number of detected HIPs varies between 2% for Mo in CHIPS to 15% for C and Au in FLUKA.
@@
In data, this issue was addressed by introducing an additional cut on the number of halo pixels (explained in Sec. 3), extracted from the cosmic data in a similar way as the energy cuts. This way, blob-like and heavy tracks produced by protons, were distinguished among the clusters initially tagged as MIPs with the energy cuts. The efficiency of the cut on the halo pixels could not be assessed in simulations, since the halo is an ASIC-specific effect present only in the measured data. However, the extracted HIPs after the implementation of this cut in data amount to 10% of the total number of HIPs in C and Mo, and to 12% in gold. These values are well within the limits of the model’s predictions for protons incorrectly tagged as MIPs. In Fig. LABEL:fig:hitmaps_tpx3a, the hitmap shows clusters from 280 annihilation events that were classified as HIPs, while fig. LABEL:fig:hitmaps_tpx3b shows those discarded as MIPs.

### v2

@@
Abstract
@@
The mechanism of antiproton-nucleus annihilation at rest is not fully understood, despite substantial previous experimental and theoretical work. In this study we used slow extracted antiprotons from the ASACUSA apparatus at CERN to measure the charged particle multiplicities and their energy deposits from antiproton annihilations at rest on three different nuclei: carbon, molybdenum and gold. The results are compared with predictions from different models in the simulation tools Geant4 and FLUKA. A model that accounts for all the observed features is still missing, as well as measurements at low energies, to validate such models.
keywords:
@@
antiproton annihilation, nucleus, Monte Carlo simulations, annihilation products, low energy, nuclear reaction1 Introduction
@@
Antiproton () annihilation on a nucleus (A) is a particular nuclear reaction in which part of the released energy is transferred to the nucleus with reduced linear and angular momentum, compared to collisions with protons or heavy ions. Consequently, the decay mechanism of nuclei heated by stopped and energetic antiprotons allows the observation of some effects characteristic of heavy-ion reactions, such as multifragment disintegration, without nuclear compression [1, 2, 3]. Our understanding of the mechanism of antiproton annihilation in nuclear matter is based on observables such as multiplicity, energy and angular distributions of secondary particles. Various models, including the Intranuclear Cascade (INC) model [4, 5], the statistical multifragmentation model [6] or the Lanzhou quantum molecular-dynamics approach [7] describe the nuclear dynamics and decay mechanism induced by antiprotons with varying degrees of success. A great amount of complexity in the antiproton-nucleus annihilation comes from the rich variety of possible reaction channels triggered by the primary annihilation mesons.
In addition, the antiproton annihilation process is still not fully understood at a quark level and the microscopic models usually include phenomenological parameters that have been tuned to reproduce early and incomplete sets of data. The quark rearrangement model explains the correct order of magnitude for the branching ratios of p annihilation at rest into three mesons. However, to account for two-body modes, kaon production and more precise predictions of branching ratios, other contributions, such as from quark-antiquark annihilation have to be included. A phenomenology was developed to try to extract from the data the relative importance of the various types of diagrams, but there is no consensus as to which model prevails [8].
@@
Annihilation at rest most of the time occurs on a single nucleon in the peripheral nuclear region, where the nucleon density is less than 10% of the central density [5]. For annihilation, the total energy of 1877 MeV is converted into an average of 3.00.2 charged pions () and 2.00.2 neutral pions (), with an average kinetic energy of 230 MeV [9, 8]. Kaons and mesons are produced in about 6% and 7% of all annihilations, respectively. When the annihilation occurs on a neutron, on average 1.070.04 and 2.070.04 are produced [10]. Absorption or scattering (final state interactions, FSI) of the emitted mesons may follow, with or without break-up of the nucleus, which depends upon the solid angle under which the remaining nucleons are seen from the annihilation point [11], and which may modify the detected meson distribution [12]. While many annihilation channels have been measured for antiproton-nucleon (N) annihilation at rest, for A, the branching ratios have been hitherto reported only for and [13, 10, 14], and the total prong (charged particle) multiplicity has been measured for a very limited number of nuclei [15]. Furthermore, even for annihilations the current models do not perform satisfactorily in describing all the aspects of the measured data.
@@
The bulk of data originates from LEAR (Low Energy Antiproton Ring) experiments, taken about three decades ago by degrading 20 MeV antiprotons in a number of moderators, and ultimately stopping them in thick targets. In these measurements, only light prongs were detected (helium and lithium ions being the heaviest), in certain energy intervals and within narrow solid angles [3, 16, 1, 17]. Features such as particle yields and momentum spectra of the emitted pions, protons (p), deuterons (d) and tritons (t) with energies above 10 MeV were reported, as well as the energy dissipation (energy transfer to the nucleons and mean excitation energies [3]). The residual nuclei have been investigated through measurements of their radioactivity after irradiation, i.e. of their gamma spectra [18, 19, 20, 21, 22, 23], and the neutron density distributions were deduced through nuclear spectroscopy analysis of the antiproton annihilation residues [24]. The experimental results have been most extensively compared to the INC, which models the A annihilation through the antiproton annihilation on a single nucleon, generating a number of pions which then cascade through the nucleus in a sequence of interactions with the rest of the nucleons. Good agreement with data was found on the absolute pion and proton yields, however the momentum distribution of pions and protons emerging from the annihilation were described with less accuracy [25]. For the residue distribution, the evaporation model provides a satisfactory description [26].
@@
Besides its relevance for nuclear physics studies, the antiproton-nucleus annihilation at rest is one of the key processes for experiments at the Antiproton Decelerator (AD) at CERN, that study antimatter in laboratory, detecting it through annihilation [27, 28, 29, 30, 31, 32]. Monte Carlo simulations are performed to assess the efficiency of tagging antihydrogen events [33], relying on the physics assumptions implemented either in Geant4, such as the CHiral Invariant Phase Space event generator (CHIPS) [34, 35, 36], Fritiof-Precompound (FTFP) [37, 38, 39] and the newly added Intranuclear Cascade de Liège (INCL) model [40], or in FLUKA [41, 42, 43]. The low energy annihilation models used in these software packages are, however, based on hadronic high energy interactions (CHIPS, FTFP) or were developed for medical physics applications (FLUKA), and were extrapolated to low energies in spite of the fact that the low-energy annihilation mechanism is still not well understood [44]. Recent studies have been performed using 5 MeV pulsed antiproton beam from the AD, degrading it to energies of 100 keV in a broad distribution. Antiprotons annihilating at rest in Al, Cu, Ag, Au and Si were measured with emulsion and silicon detectors, as reported in [45, 46, 47]. The results revealed significant discrepancies in the prong multiplicity between measured data and the Geant4 models, with differences ranging from 30% to a factor of 4.
In this work, we present new measurements of A annihilation at rest using a slow extracted, monoenergetic beam of 150 eV, accelerated to 1.15 keV by the electric potential applied on the solid target. The antiprotons annihilated on thin targets, providing a crucial comparison with the models in which the stopping effects and production of secondaries due to the particle propagation can be neglected. In addition, this allows the emergence and detection of the heavy nuclei produced in the annihilation. The presented work initiates an extended experimental study at ELENA [48], along with state-of-the art detector technologies, to study in detail the production of various charged particles in antiproton induced reactions in different nuclei (see Sec. 5). The ultimate goal is to provide a deeper understanding of antiproton-nucleus annihilation through quantitative and qualitative analyses of FSIs and their and their dependence on the atomic number.
@@ 2 Experimental setup
The antiprotons were annihilated on three different targets: 2 thick diamond-like-carbon (DLC) and molybdenum foils, and a gold foil of 1 thickness, biased at 1 kV. The stopping range of 1.15 keV antiprotons, according to ref. [55], is 46 nm for carbon and 57 nm for gold. The beam was centered and focused on the target with an Einzel lens (see Fig. 1), where both the entrance and exit electrodes have additional diagonal cuts. Consisting of five electrodes, this lens allows for two-dimensional steering perpendicularly to the beam axis.
The on-target beam size was determined by first measuring the antiproton beam on the bismuth germanium oxide (BGO) crystal detector used for the antihydrogen experiment [56], and by using the results of 30 mm FWHM in x and y direction [57] as an input to simulations. As the foils then replaced the BGO for the current measurements, the beam size on the target foils was estimated from the measured size of the beam, adding, in the Geant4 simulations, the 1 kV electric potential applied on the foil for further focusing. The outcome of this combined work is shown in Fig. 2a and 2b, where the x-y beam profile and its horizontal (x) and vertical (y) projections are shown, respectively. A beam size of 10.5 mm in x and y (FWHM) was obtained, with 92.4 of the extracted antiprotons annihilating on the 22 target, and the remainder on the mechanical support. These results were also confirmed with SIMION [58] simulations, starting from the same initial conditions and including the voltage bias on the foil.
@@ 2.1 Detection and identification of charged pions and heavy fragments
The charged annihilation products were simultaneously detected by two detectors, a cylindrical hodoscope surrounding the target and a pixel detector placed 1 cm behind it (Timepix3 quad), as shown in Fig. 3. Pions and heavier fragments (p, d, t, He, Li, etc.) in the forward direction were impinging on the 22 array of Timepix3, with a total area of 2.82.8 . The hybrid pixel detector (HPD) readout chip, containing 256256 pixel channels of 5555 each, provides a simultaneous information of the time-of-arrival (ToA) and the time-over-threshold (ToT) for each pixel, with a nominal time resolution of 1.56 ns. The choice of the sensor depends on the particular application, and, as silicon is well suited for detection of charged particles, the quad was coupled to a 500 thick silicon sensor. The detector is able to detect mixed radiation fields, composed of various types of particles, and has a long history of a variety of applications [59, 60, 61, 62, 63]. Its wide dynamic spectral range allows for detection of all charged particles depositing energy above the threshold. In this experiment, the detection threshold was set to of collected charge per pixel, corresponding to a minimum required energy of 3.6 keV for X rays and low-energy gamma rays, tens of keV for electrons, hundreds of keV for protons and MeV for heavy ions [64]. The higher threshold for heavier particles originates from the aluminum electrode, approximately 500 nm in thickness, that is deposited on top of the silicon sensor for biasing purposes. For details about the ASIC and the readout system see ref. [65, 66, 67].
@@
The two detectors were time synchronised, and the data acquisition coincided with the start of the slow extraction cycle. Every time a hit was detected in both inner and outer bar layers of the hodoscope, a trigger was issued and this event was time-stamped in the Timepix3 data stream, where every hit in this detector was also recorded. The time distribution of these events (triggers) for one extraction cycle is shown in Fig. 4a, revealing the 20 s duration of the antiproton beam extracted from MUSASHI within the 140 s long cycle. The double peak structure is related to the DC extraction, and possible sources are the imperfect cooling of the antiprotons, or a mismatch to the space charge potential when the well depth was slowly decreased during the extraction. About 1,000 annihilation events per slow extraction cycle were detected, resulting in a total of 100,000 annihilation events for each target foil. The triggers outside the slow extraction window are produced by cosmic rays or, sporadically, from secondaries of other origin in the AD experimental hall crossing the hodoscope. The two small peaks in the Timepix3 distribution are from two cosmic particles that traversed both the hodoscope and the Timepix3 detectors. The data analysis was performed only within the slow extraction time window.
@@
The hits in the Timepix3 were assigned to a particular event (trigger) according to their ToA. The interval between two successive triggers ranged from 10 ms to a few 10 ms, as illustrated in Fig. 4b. The observed cutoff at 10 ms stems from the hodoscope data acquisition limit of 100 Hz. The nominal cosmic background was measured to be 1.3 Hz. When combined with the duration of one extraction (20 s) it gives a total of 26 cosmics events, which is only 2.6% of the average number of annihilation events recorded per run. For the Timepix3 the cosmic ray background is fully negligible.
@@ 2.1.2 Charged pions and heavier fragments in Timepix3
In the pixel detector, charged particles were detected in 25% of the full solid angle. Based on the ToT proportionality to the amount of charge collected in each electrode, a per pixel energy calibration was carried out using test pulses. The deposited energy was extracted from the known value of the test pulse capacitor, assuming an energy of 3.6 eV for the creation of one electron-hole pair in silicon. The calibration results were validated with data from a radioactive source and cosmic rays. In Fig. 5a, the lower end of the spectrum, converted to energy units, is presented with a Gaussian fit to the characteristic 59.5 keV peak, determined to be 59.32.6 keV.
@@
The energy deposit and morphology of the pixelated track that a charged particle generates in the Timepix3 (also referred to as pixel cluster) are influenced by factors such as the particle type, its kinetic energy, and the incidence angle. The graph in Fig. 5b illustrates the distribution for cosmic rays, both in measurements and simulations, exclusively for clusters with high (0.8) eccentricity, a measure of the linearity of a particle track. These elongated tracks correspond to cosmic rays that cross the Timepix3 at large angles with respect to its normal. The distribution is modeled with a Landau-Gaussian fit and compared to Geant4 simulations. The most probable value of 1.4 agrees with the typical value characteristic for MIPs in silicon [70].
@@
Due to the geometry of the set-up (see Fig. 3), the annihilation prongs hit the pixel detector only at acute angles with respect to its normal. The maximum angle is 55 for prongs emerging from the foil’s center, and up to 67 for prongs originating from annihilations close to the edges of the target. Most of the particles impinge on the Timepix3 at angles smaller than these, which results in the formation of mostly blob-like clusters and relatively short tracks, with a maximum length of less than 1.5 mm. Combined with the variety of particles produced in the annihilation and their diverse ranges of energies and incident angles, this allowed only a broad classification of heavily ionizing particles (HIPs), as opposed to MIPs. In this work, protons and all heavier nuclear fragments are classified as HIPs, regardless of their energy. Ongoing efforts from a number of users of Timepix ASICs are aimed at developing tools for more precise particle identification in mixed radiation fields [64].
@@
In this study, the differentiation of MIPs from HIPs in the Timepix3 was achieved by studying the cluster morphology and energy deposits from a cosmics data sample. Since the distribution of the energy deposit from a cosmic, or a MIP in silicon partly overlaps with that of protons with energies of 100 MeV, other criteria, more specific to the cosmics particles were required to distinguish MIPs in the annihilation data. Hence, as a first step, a double cut was derived from the analysis of cosmic data on the highest pixel energy () within the cluster and the mean pixel energy (). The energy thresholds were extracted from the versus distribution depicted in Fig. 6a, by calculating its cumulative distribution function (CDF), shown in Fig. 6b. The thresholds were set to account for 99 of the cosmic rays data sample, leading to values of 55 keV and 164 keV for and respectively. The remaining clusters underwent further evaluation based on their morphology and were categorized as small blobs, large blobs, or heavy tracks before being identified as HIPs (see Fig. 8a). The efficiency of removing MIPs from the annihilation clusters using only energy cuts from cosmics was estimated, from simulations, to be 97. However, it was also revealed that some high-energy protons can be mistakenly identified as MIPs after applying these cuts, as shown in Fig. 7. The comparison plots show the deposited energy by charged pions (solid line) and by MIPs tagged after applying the energy cuts obtained from simulated cosmic rays (dashed lines) for two Geant4 models ((a) FTF and (b) CHIPS), for all three targets. Additionally, the difference between the two energy depositions is presented at the bottom. The excess counts are attributed to the aforementioned protons, the quantity of which depends on the specific simulation model. The percentage of these protons incorrectly tagged as MIPs relative to the total number of detected HIPs varies between 2% for Mo in CHIPS to 15% for C and Au in FLUKA.
@@
In data, this issue was addressed by introducing an additional cut on the number of halo pixels (explained in Sec. 3), extracted from the cosmic data in a similar way as the energy cuts. This way, blob-like and heavy tracks produced by protons, were distinguished among the clusters initially tagged as MIPs with the energy cuts. The efficiency of the cut on the halo pixels could not be assessed in simulations, since the halo is an ASIC-specific effect present only in the measured data. However, the extracted HIPs after the implementation of this cut in data amount to 10% of the total number of HIPs in C and Mo, and to 12% in gold. These values are well within the limits of the model’s predictions for protons incorrectly tagged as MIPs. In Fig. 8a, the hitmap shows clusters from 280 annihilation events that were classified as HIPs, while Fig. 8b shows those discarded as MIPs.